### v1

@@ Abstract.
This paper presents the first rotating leader state machine replication (SMR) protocol that allows transactions to be confirmed in just a single round-trip time in the Byzantine fault tolerance (BFT) setting. Based on minimal alterations to the Internet Computer Consensus (ICC) protocol and with negligible communication overhead, we introduce a novel dual mode mechanism that enables optimal block finalization latency in the fast path. Crucially, the modes of operation are integrated, such that even if the fast path is not effective, no penalties are incurred. Moreover, our algorithm maintains the core attributes of the original ICC protocol, including optimistic responsiveness and rotating leaders without the necessity for a view-change protocol.
We prove the correctness of our Fast Internet Computer Consensus (FICC) protocol and provide an open-source implementation of it. Both the FICC and original ICC protocol are compared in a globally distributed wide-area network. Our evaluation reveals that the FICC protocol achieves reduced latency compared to the ICC protocol, without requiring additional security assumptions. Furthermore, by increasing the number of replicas to $n=5f+1$, we exhibit that latency improvements close to the theoretical maximum of 33% are attainable. We conclude by highlighting the network topology as a significant factor in evaluating and comparing the latency of consensus algorithms.
state machine replication, consensus, byzantine fault tolerance, blockchain, fast path
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Security and privacy Distributed systems securityccs: Applied computing Secure online transactions
@@ 1. Introduction
Popularized by Bitcoin (Nakamoto, 2009), Ethereum (Buterin, 2013) and other digital currencies, protocols that tolerate an open set of participants have emerged. This has not lessened interest in the permissioned setting, however, as their approaches are often based on advancements from permissioned counterparts (see Ethereum’s Casper (Buterin and Griffith, 2017) and Algorand (Gilad et al., 2017) for instance).
The practicality of these systems has been shown in the context of finance and digital art (Uniswap, 2023; Compound, 2023; MakerDAO, 2023; OpenSea, 2023), but other promising applications such as file storage, social media, governance, and democracy have emerged (Filecoin, 2023; dscvr, 2023; Tally, 2023; Worldcoin, 2023). In essence, BFT protocols can be applied whenever multiple parties desire to perform some computation, without trusting any set of participants fully.
@@
These decentralized systems are not unlike a decentralized world computer, enabling anyone to run Turing-complete programs. Various such world computers exist today (Buterin, 2013; Dfinity, 2023c; Filecoin, 2023; Blackshear et al., 2023), with varying trade-offs regarding performance, security, and inclusiveness. At each system’s core, lays an ordering protocol assuring that resource accesses are consistent across all participants (called replicas), thus guaranteeing deterministic and secure execution of the computation.
@@
Byzantine agreement (BA) protocols typically provide consensus on a single decision, while state machine replication protocols (SMR) focus on making efficient use of this costly primitive. In practice, the most widely used SMR protocols (such as Bitcoin (Nakamoto, 2009), Ethereum (Buterin, 2013), and Algorand (Gilad et al., 2017)) make use of an elected leader to propose a batch of transactions, called a block. A total order across leaders is then obtained by chaining blocks.
@@
There are a few considerations that are especially crucial to such leader-based protocols. Firstly, leaders can misbehave and propose conflicting blocks. While this can be detected, the resulting change of the leader (called view-change in the literature) leads to notoriously high complexity. Honest, but slow leaders are just as problematic, as they cannot easily be called out, but can stall the effective throughput nonetheless (Clement et al., 2009).
Moreover, relying on a single leader leads to uneven load, both in terms of computation and communication (Crain et al., 2018). Finally, leaders can sometimes extract value from their control of the block creation. On the Ethereum blockchain for example, miner-extractable value (MEV) is a major source of income for replicas and leads to potentially dangerous incentives to deviate from honest behavior (Daian et al., 2020).
These issues can at least partially be resolved by employing rotating or random leaders (Camenisch et al., 2022), which is the class of protocols we aim to improve upon in this work.
Figure 1. The FICC protocol can terminate after just two communication steps. Existing rotating leader SMR protocols require at least three communication steps.
@@
Specifically, we address what we believe to be a major roadblock in the adoption of decentralized computers: high latency. In contrast to privacy and security, latency is at the forefront of the user experience. While centralized and trusted applications (such as credit card payments or social media) provide almost instantaneous services, their Byzantine fault tolerant counterpart typically comes at the cost of a noticeable delay, in some cases in the order of minutes or even hours (such as in Bitcoin). Furthermore, contrary to other metrics such as throughput, latency cannot be readily improved by hardware upgrades.
@@
In this work, we present the first rotating leader SMR protocol that allows transactions to be confirmed in just a single round-trip time. This is optimal, as to achieve fault tolerance the block must 1) reach a fraction of replicas, that have to 2) respond to acquiesce to the processing. This is visualized in fig. 1.
@@
We are able to achieve this by uniting two lines of research. The first deals with the theoretical bounds of latency, and has established optimally fast single-shot protocols in the good case (Kursawe, 2002; Martin and Alvisi, 2006; Aublin et al., 2015; Kuznetsov et al., 2021; Abraham et al., 2021b). The second, more well-known line of research, provides a class of consensus protocols that solve state machine replication, in an efficient and fair way by employing rotating leaders. Prominent examples include HotStuff, Tendermint, Casper, Algorand and Bullshark (Yin et al., 2019; Buchman, 2016; Buterin and Griffith, 2017; Gilad et al., 2017; Spiegelman et al., 2022).
@@
As Chan and Pass point out (Chan and Pass, 2023), rotating leader protocols can be differentiated by either favoring proposal confirmation time, meaning block finalization latency, or block creation time, corresponding to the delay between the blocks and thus the associated throughput. The Internet Computer Consensus family of protocols (Camenisch et al., 2022) excels in both of these metrics. The protocol powers the Internet Computer (IC) project (Dfinity, 2023c): a decentralized platform enabling developers to deploy dApps that run fully on the blockchain. The system is composed of multiple permissioned sets of replicas running datacenter-grade infrastructure managed by a decentralized autonomous organization.
Exhibiting around one to two seconds of block finalization latency, this deployment of ICC is among the quickest consensus algorithms in the blockchain space. Yet, it is arguably not fast enough to compete with the Web services hosted on centralized server infrastructures.
@@
Our contribution. We present the Fast Internet Computer Consensus (FICC) protocol that achieves optimal one-shot consensus latency while preserving the state-of-the-art characteristics of ICC.
@@
    FICC is the first SMR protocol with rotating leaders supporting two-step chain growth and two-step finalization under partial synchrony. We prove the safety and liveness of FICC for $n\geq 3f+2p+1$, where $n$ is the total number of replicas in the network, and $f$ is the maximum number of Byzantine replicas tolerated. Additionally, $p$ is a parameter that can be set freely ($0\leq p\leq f$), and determines the effectiveness of the fast path. We prove that as long as no more than $p$ replicas are unresponsive, FICC block latency is just two network delays.
@@
    We provide an experimental FICC implementation in Rust, forked from the Internet Computer codebase.
@@
    We measure the proposal finalization time in a wide area network spanning the globe. We empirically show that FICC improves proposal finalization time compared to the reference ICC implementation without any relaxation of the security guarantees.
@@ 2. Related work
Fast Consensus. Our algorithm relies on a long line of established work on reaching consensus in two communication steps (or one round trip time), also referred to as fast or early-stopping consensus (Brasileiro et al., 2001; Kursawe, 2002; Song and van Renesse, 2008; Martin and Alvisi, 2006; Guerraoui and Vukolić, 2007). As shown by Brasileiro et al. (Brasileiro et al., 2001) in the crash-fault model, it is possible to terminate early if “enough” acknowledgments are received. A year later, a fast path was presented by Kursawe (Kursawe, 2002), in which a two-step fast path is paired with a subprotocol in the slow path. Upon the expiration of a timer, a fallback subprotocol is used to ensure liveness.
@@
Song et al. describe Bosco (Song and van Renesse, 2008), an algorithm that provides a two-step fast path to any underlying consensus algorithm. Their fast path does not rely on synchrony assumptions, and, assuming $n>5f$, is triggered when all servers are in pre-agreement. Assuming $n>7f$, the fast path is triggered when all honest servers are in pre-agreement instead. Using similar building blocks it was later shown that consensus can be performed “on demand”, i.e., the fallback is only used when the fast path does not succeed (Sliwinski et al., 2022). As fig. 2 visualizes, these approaches come at the cost of a longer recovery during the fallback on the slow path though, as the slow path can only be started once the fast path round is over. In this work, we show that both the (single) fast path round and the first slow path round can happen simultaneously (with only constant message overhead).
@@
Figure 2. The different approaches of fast path protocols. In many protocols, the slow path starts after the fast path fails. In FICC, the fast path is integrated into the slow path.
@@
Earlier, Martin and Alvisi introduce Fast Byzantine Paxos (FaB Paxos) (Martin and Alvisi, 2006), a family of protocols parameterized not only by $f$ but also $p$ ($0\leq p\leq f$), the number of (non-leader) replicas that are not needed for the fast path to succeed. Their algorithms provide Byzantine agreement, when $n\geq 3f+2p+1$, where the fast path triggers as long as $n-p$ servers behave honestly and are in pre-agreement. We use the same security model and guarantees in this work, however, we do not require progress certificates to guarantee safety. Additionally, instead of relying on a variant of PBFT in the slow path, we integrate our protocol with ICC, sidestepping the complexity introduced by view changes.
The lower bound by Martin and Alvisi (Martin and Alvisi, 2006) has recently been improved to $n\geq max(3f+2p-1,3f+1)$ by Kuznetsov et al. (Kuznetsov et al., 2021) and Abraham et al. (Abraham et al., 2021b). This is applicable when servers and clients are identical sets. Their insight to observe and discard actions by provably misbehaving leaders does not trivially apply to our use case, as we propose an SMR algorithm instead of a single-shot consensus instance. A more rigorous analysis is left for future work.
@@
Fast SMR. Ideas from the FaB protocol were borrowed by the Zyzzyva protocol (Kotla et al., 2007) as well, which turns the single-shot FaB consensus into an SMR protocol. Improving upon Zyzzyva, Aliph (Aublin et al., 2015) reduces the latency to two steps. These protocols do not allow running both the fast path and a slow path concurrently though, and thus suffer from switching costs (see fig. 2). Zyzzyva safety issues were pointed out by (Abraham et al., 2017) and lead to the development of Zelma (Abraham et al., 2018), combining FaB and Zyzzyva’s benefits.
@@
Zelma is also at the core of the SBFT protocol (Gueta et al., 2019). SBFT is the first protocol running both pessimistic and optimistic paths simultaneously in a dual mode. However, it requires more communication steps when compared to FICC. Indeed, this approach lends itself to ICC exceptionally well, as the protocol tolerates arbitrarily delaying the decision on a value until (in a phase of synchrony) a successful consensus instance retroactively finalizes previous instances. A large body of work has based itself on the 3-phase (6-step) HotStuff (Yin et al., 2019) protocol and presented improvements upon it. Jolteon and Ditto (Gelashvili et al., 2022) have improved the latency while adding an asynchronous fallback to enhance its performance during epoch synchronization (see table 1).
@@
Most recently Malkhi and Nayak propose a 2-phase (4-step) protocol and a simpler fast path (Malkhi and Nayak, 2023). It is our understanding that these protocols reach no lower than $3\delta$ proposer latency in the fast path.
Rotating Leader BFT protocols. Mir-BFT (Stathakopoulou et al., 2019) and its follow-up work (Stathakopoulou et al., 2022) improve upon sequential-leader approaches, by running PBFT instances on a set of leaders. When a leader is slow, it is replaced. Compared to (F)ICC protocols, it suffers from the complexity of view-changes. Other SMR protocols are designed to perform well under attack (Avarikioti et al., 2023). However, in the permissioned blockchain setting it is often assumed that during periods of synchrony, misbehavior can be punished, either by exclusion or slashing (Buterin and Griffith, 2017). As such we focus on the latency in the optimistic case and satisfy the same pessimistic liveness as ICC, which was shown to be adequate in (Chan and Pass, 2023). Recent concurrent work by Chan and Pass (Chan and Pass, 2023) provides a theoretical framework for this class of algorithms, and table 1 is partially inspired by it. Their proposed protocol called Simplex falls into the same category as the ICC and FICC algorithms but improves pessimistic liveness by only allowing a single leader per round. We believe that a technique similar to the one shown in this work could be applied to bring our fast path to Simplex too. In (Abraham et al., 2021a) Abraham et al. prove that $2\Delta$ is the upper and lower bound for the good-case latency of rotating leader BFT protocols, in the synchronous setting. In this work, we provide a matching upper bound in the partially synchronous setting.
DAG-based BFT protocols. Another line of work proposes to solve consensus efficiently by using DAG-based protocols (Keidar et al., 2021; Spiegelman et al., 2022; Babel et al., 2023; Malkhi et al., 2023).
One main idea is to disconnect transactions broadcasting from finalization (Danezis et al., 2022), and another is to allow the finalization of blocks outside the main blockchain by causally referencing them (Lewenberg et al., 2015; Keidar et al., 2021). In Bullshark (Spiegelman et al., 2022), one of the most recent protocols, all replicas are parallel leaders and broadcast one batch of transactions in each round using Byzantine Consistent Broadcast. While doing so, they include references to at least $n-f$ blocks from the previous round. A more recent protocol called BBCA-Chain (Malkhi et al., 2023) reduces Bullshark’s latency by removing the need for specialized block layers, and instead broadcasts blocks using a primitive called Byzantine Broadcast with Complete-Adopt (BBCA). (F)ICC does not allow causally referencing non-leader blocks to finalize them and thus doesn’t benefit from the same throughput advantages. However, we believe ICC is a great fit to describe our fast path mechanism, which might be applicable to other DAG protocols too.
In table 1 an overview of state-of-the-art protocols is shown. For both the proposer latency and the block creation latency, we have included the required number of replicas that have to respond in order to proceed in the optimistic case (synchronous round with a block proposed by an honest leader). While irrelevant in the calculation of the theoretical latency bounds, we found that small changes in these requirements could have large effects in practice, especially on global-scale deployments as they are seen today. (Intuitively, it is not uncommon for a few outlier nodes to have a higher latency. If progress can be made safely without them, a large performance increase can be experienced.)
Block Finalization
@@
Casper FFG (Buterin and Griffith, 2017) $O(\Delta)$ $\ {{\lceil\frac{2n}{3}\rceil}}$ $O(\Delta)$ N/A *
@@
PaLa (Chan et al., 2018) $4\delta$ ${\lceil\frac{2n}{3}\rceil}$ $2\delta$ ${\lceil\frac{2n}{3}\rceil}$ $13f+$ 1
Zelma (Abraham et al., 2018) $2\delta$ $n-p$ $2\delta$ $n-f-p$ $13f+2p+$ 1
SBFT (Gueta et al., 2019) $3\delta$ $13f+p+$ 1 $3\delta$ $13f+p+$ 1 $13f+2p+$ 1
Bullshark (Spiegelman et al., 2022) $4\delta$
$n-f$ with
$f+1$ overlap
$2\delta$
$n-f$ (odd rounds)
$f+1$ / $12f+$ 1 (even rounds)
$13f+$ 1
@@
ICC (Camenisch et al., 2022)
Simplex (Chan and Pass, 2023)
$3\delta$ $n-f$ $2\delta$ $n-f$ $13f+$ 1
FICC $2\delta$ $n-p$ $2\delta$ ${\lfloor\frac{n+f}{2}\rfloor}+1$ $13f+2p+$ 1
@@
Table 1. Popular and Fast State Machine Protocols. *Non-equivocation is enforced by slashing. † To the best of our knowledge no pipelining is specified. $\Delta$ denotes the message delivery time upper bound, while $\delta$ is the true message delivery time.
@@
Other aspects. Many works such as (Malkhi and Nayak, 2023; Camenisch et al., 2022; Gueta et al., 2019) also focus on the message complexity of the fast path and of the view change. We do describe how BLS signature aggregation (Boneh et al., 2018) can be used for FICC, but leave further advancements to future work. Message complexity and performance do not always go hand in hand in practice (Camenisch et al., 2022) and oftentimes probabilistic broadcast through gossiping is preferred (Protocol Labs, 2023).

### v2

@@ Abstract.
This paper presents Banyan, the first rotating leader state machine replication (SMR) protocol that allows transactions to be confirmed in just a single round-trip time in the Byzantine fault tolerance (BFT) setting. Based on minimal alterations to the Internet Computer Consensus (ICC) protocol and with negligible communication overhead, we introduce a novel dual mode mechanism that enables optimal block finalization latency in the fast path. Crucially, the modes of operation are integrated, such that even if the fast path is not effective, no penalties are incurred. Moreover, our algorithm maintains the core attributes of the ICC protocol it is based on, including optimistic responsiveness and rotating leaders without the necessity for a view-change protocol.
We prove the correctness of our protocol and provide an open-source implementation of it. Banyan is compared to its predecessor ICC, as well as other well known BFT protocols, in a globally distributed wide-area network. Our evaluation reveals that Banyan reduces latency by up to 30% compared to state-of-the-art protocols, without requiring additional security assumptions.
Consensus, Blockchain, Byzantine fault tolerance, Fast Path, State Machine Replication
ccs: Computer systems organization Distributed architecturesccs: Security and privacy Distributed systems security
@@ 1. Introduction
This core primitive has enabled the creation of decentralized protocols that are akin to a world computer, enabling anyone to run Turing-complete programs. Various such world computers exist today (Buterin, 2013; Dfinity, 2024b; Filecoin, 2024; Blackshear et al., 2023), with varying trade-offs regarding performance, security, and inclusiveness. At each system’s core, lays an ordering protocol assuring that resource accesses are consistent across all participants (called replicas), thus guaranteeing deterministic and secure execution of the computation.
@@
Byzantine agreement (BA) protocols typically provide consensus on a single decision, while state machine replication protocols (SMR) focus on making efficient use of this costly primitive. In practice, the most widely used consensus protocols (such as Bitcoin (Nakamoto, 2009), Ethereum (Buterin, 2013), and Algorand (Gilad et al., 2017)) make use of an elected leader to propose a batch of transactions, called a block. A total order across leaders is then obtained by chaining blocks.
@@
There are a few considerations that are especially crucial to such leader-based protocols. Firstly, leaders can misbehave and propose conflicting blocks. While this can be detected, the resulting change of the leader (called view-change in the literature) leads to notoriously high complexity. Honest, but slow leaders are just as problematic, as they cannot easily be called out, but can stall the effective throughput nonetheless (Clement et al., 2009). Moreover, relying on a single leader leads to uneven load, both in terms of computation and communication (Crain et al., 2018), and might facilitate censorship. Finally, leaders sometimes extract value from their control of the block creation. On the Ethereum blockchain, for example, miner-extractable value (MEV) is a major source of income for replicas and leads to potentially dangerous incentives to deviate from honest behavior (Daian et al., 2020). These issues can at least partially be resolved by employing rotating or random leaders (Camenisch et al., 2022), which is the class of protocols we aim to improve upon in this work.
@@
Specifically, we address what we believe to be a major roadblock in the adoption of decentralized computers: high latency. In contrast to privacy and security, latency is at the forefront of the user experience. While centralized and trusted applications (such as credit card payments or social media) provide almost instantaneous services, their Byzantine fault-tolerant counterpart typically comes at the cost of a noticeable delay, in some cases in the order of minutes or even hours (such as in Bitcoin). Furthermore, contrary to other metrics such as throughput, latency cannot be readily improved by hardware upgrades.
Figure 1. Banyan can terminate after just two communication steps. Existing rotating leader BFT protocols require at least three communication steps.
@@
In this work, we present the first rotating leader SMR protocol that allows transactions to be confirmed in just a single round-trip time. This is optimal, as to achieve fault tolerance the block must 1) reach a fraction of replicas, that have to 2) respond to acquiesce to the processing (Song and van Renesse, 2008). This is visualized in Figure 1.
@@
We are able to achieve this by uniting two lines of research. The first deals with the theoretical bounds of latency, and has established optimally fast single-shot protocols in the good case (Kursawe, 2002; Martin and Alvisi, 2006; Aublin et al., 2015; Kuznetsov et al., 2021; Abraham et al., 2021b). The second, more well-known line of research, provides a class of consensus protocols that solve state machine replication, in an efficient and fair way, by employing rotating or random leaders. Prominent examples include HotStuff, Tendermint, Casper, Algorand and Bullshark (Yin et al., 2019; Buchman, 2016; Buterin and Griffith, 2017; Gilad et al., 2017; Spiegelman et al., 2022).
@@
As Chan and Pass point out (Chan and Pass, 2023), rotating leader protocols can be differentiated by either favoring proposal confirmation time, meaning block finalization latency, or block creation time, corresponding to the delay between the blocks and thus the associated throughput. The Internet Computer Consensus family of protocols (Camenisch et al., 2022) excels in both of these metrics.
@@
Our contribution. We present the Banyan protocol that achieves optimal consensus latency while preserving the state-of-the-art characteristics of ICC.
@@
    Banyan is the first SMR protocol with rotating leaders supporting two-step chain growth and two-step finalization under partial synchrony. We prove the safety and liveness of Banyan for $n\geq 3f+2p^{*}-1$, where $n$ is the total number of replicas in the network, and $f$ is the maximum number of Byzantine replicas tolerated. Additionally, $p^{*}\geq 1$ is a parameter that can be set freely ($p^{*}\leq f$), and determines the effectiveness of the fast path. We prove that as long as no more than $p^{*}$ replicas are unresponsive, Banyan finalization latency is just two network delays.
@@
    We provide a proof-of-concept Banyan implementation written in Golang, built on top of the Bamboo BFT framework (Gai et al., 2021).
@@
    We measure the proposal finalization time in three different wide area networks. We empirically demonstrate that, compared to ICC, HotStuff, and Streamlet, Banyan achieves the fastest proposal finalization time, improving by up to 30% over the runner-up. Further, we show that Banyan does not suffer from increased variance in its latency, and behaves as ICC under crash-faults.
@@ 2. Related Work
Fast Consensus. Our algorithm relies on a long line of established work on reaching consensus in two communication steps (or one round trip time), also referred to as fast or early-stopping consensus (Brasileiro et al., 2001; Kursawe, 2002; Friedman et al., 2005; Song and van Renesse, 2008; Martin and Alvisi, 2006; Guerraoui and Vukolić, 2007). As shown by Brasileiro et al. (Brasileiro et al., 2001) in the crash-fault model, it is possible to terminate early if “enough” acknowledgments are received. A year later, a fast path was presented by Kursawe (Kursawe, 2002), in which a two-step fast path is paired with a subprotocol in the slow path. Upon the expiration of a timer, a fallback subprotocol is used to ensure liveness.
@@
Song et al. describe Bosco (Song and van Renesse, 2008), an algorithm that provides a two-step fast path to any underlying consensus algorithm. Their fast path does not rely on synchrony assumptions, and, assuming $n>5f$, is triggered when all servers are in pre-agreement, i.e., propose the same value. Assuming $n>7f$, the fast path is triggered when all honest servers are in pre-agreement instead. Using similar building blocks it was later shown that consensus can be performed “on demand”, i.e., the fallback is only used when the fast path does not succeed (Sliwinski et al., 2022). As Figure 2 shows, these approaches come at the cost of a longer recovery during the fallback on the slow path though, as the slow path can only be started once the fast path round is over. In this work, we show that both the (single) fast path round and the first slow path round can happen simultaneously (with only constant message overhead).
@@
Figure 2. The different approaches of fast path protocols. In many protocols, the slow path starts after the fast path fails. In Banyan, the fast path is integrated into the slow path.
@@
Earlier, Martin and Alvisi introduce Fast Byzantine Paxos (FaB Paxos) (Martin and Alvisi, 2006), a family of protocols parameterized not only by $f$ but also $p$ ($0\leq p\leq f$), the number of (non-leader) replicas that are not needed for the fast path to succeed. Their algorithms provide Byzantine agreement, when $n\geq 3f+2p+1$, where the fast path triggers as long as $n-p$ servers behave honestly and are in pre-agreement. Instead of relying on a variant of PBFT in the slow path, we integrate our protocol with ICC, sidestepping the complexity introduced by view changes.
The lower bound by Martin and Alvisi (Martin and Alvisi, 2006) has recently been improved to $n\geq max(3f+2p-1,3f+1)$ by Kuznetsov et al. (Kuznetsov et al., 2021) and Abraham et al. (Abraham et al., 2021b). Their insight is that misbehaving (i.e., equivocating) leaders can be reliably detected, and thus acceptors can wait for $n-f$ votes, excluding the malicious leader ($n-f+1$ votes in total). To the best of our knowledge, we present the first protocol that matches this lower bound without the need for a view-change protocol.
@@
Fast SMR. Ideas from the FaB protocol were borrowed by the Zyzzyva protocol (Kotla et al., 2007) as well, which turns the single-shot FaB consensus into an SMR protocol. Improving upon Zyzzyva, Aliph (Aublin et al., 2015) reduces the latency to two steps. These protocols do not allow running both the fast path and a slow path concurrently, though, and thus suffer from switching costs (see Figure 2). Zyzzyva safety issues were pointed out by (Abraham et al., 2017) and led to the development of Zelma (Abraham et al., 2018), combining FaB and Zyzzyva’s benefits.
@@
Zelma is also at the core of the SBFT protocol (Gueta et al., 2019). SBFT is the first protocol running both pessimistic and optimistic paths simultaneously in a dual mode. However, the fast path of SBFT has one more communication step than Banyan, and assumes $n\geq 3f+2p+1$. Moreover, the slow path of SBFT is not optimistically responsive, as it triggers only after a time-out (fig. 2). A large body of work has based itself on the 3-phase (6-step) HotStuff (Yin et al., 2019) protocol and presented improvements upon it (Jalalzai et al., 2020; Malkhi and Nayak, 2023; Kang et al., 2024). Jolteon and Ditto (Gelashvili et al., 2022) have improved the latency while adding an asynchronous fallback to enhance its performance during epoch synchronization (see Table 1).
@@
Most recently Malkhi and Nayak propose a 2-phase (4-step) protocol and a simpler fast path (Malkhi and Nayak, 2023). These protocols reach no lower than $3\delta$ proposer latency in the fast path.
Block
Finalization
@@
[…]
@@
Jolteon (Gelashvili et al., 2022)
$5\delta$ $12f+$ 1 $2\delta$ $12f+$ 1 $13f+$ 1
PaLa (Chan et al., 2018) $4\delta$ $12f+$ 1 $2\delta$ $12f+$ 1 $13f+$ 1
Zelma (Abraham et al., 2018) $2\delta$ $13f+p+$ 1 $2\delta$ $12f+p+$ 1 $13f+2p+$ 1
SBFT (Gueta et al., 2019) $3\delta$ $13f+p+$ 1 $3\delta^{\mathsection}$ $12f+p+$ 1 $13f+2p+$ 1
Streamlet (Chan and Shi, 2020)
$6\Delta$ $12f+$ 1 $2\Delta$ $12f+$ 1 $13f+$ 1
Bullshark (Spiegelman et al., 2022) $4\delta^{\mathparagraph}$ $12f+$ 1 $2\delta$ $12f+$ 1 $13f+$ 1
@@
ICC (Camenisch et al., 2022) / Simplex (Chan and Pass, 2023)
$3\delta$ $12f+$ 1 $2\delta$ $12f+$ 1 $13f+$ 1
Mysticeti (Babel et al., 2023)
$3\delta$ $12f+$ 1 $1\delta$ $12f+$ 1 $13f+$ 1
Banyan ${2\delta}$ $13f+p^{*}-$ 1 $2\delta$ $2f+p^{*}$ $13f+2p^{*}-$ 1
@@
Table 1. Popular and Fast State Machine Replication Protocols. To simplify comparison, we assume that the number of replicas is equal to the respective lower bound. $\Delta$ denotes the message delivery time upper bound, while $\delta$ is the true message delivery time. Non-equivocation is enforced by slashing. We consider the pipelined version of Fast HotStuff. § To the best of our knowledge no pipelining is specified for SBFT. We consider Bullshark’s best case latency (anchor blocks). For simplicity, we replace $p$ by $p^{*},p^{*}\geq 1$.
Rotating Leader BFT protocols. Many recent advances in BFT protocols rely on rotating leaders (Buchman, 2016; Buchnik and Friedman, 2019; Yin et al., 2019; Chan and Shi, 2020), as originally introduced by Veronese et al. (Veronese et al., 2009). Mir-BFT (Stathakopoulou et al., 2019) and its follow-up work (Stathakopoulou et al., 2022) for example, improve upon sequential-leader approaches, by running PBFT instances on a set of leaders. When a leader is slow, it is replaced. Other SMR protocols are designed to perform well under attack (Amir et al., 2010; Avarikioti et al., 2023). However, in the permissioned blockchain setting it is often assumed that during periods of synchrony, misbehavior can be punished, either by exclusion or slashing (Buterin and Griffith, 2017). As such, we focus on the latency in the optimistic case and satisfy the same pessimistic liveness as ICC, which was shown to be adequate in (Chan and Pass, 2023). Recent concurrent work by Chan and Pass (Chan and Pass, 2023) provides a theoretical framework for this class of algorithms, and Table 1 is partially inspired by it. Their proposed protocol called Simplex falls into the same category as the ICC and Banyan algorithms but improves pessimistic liveness by only allowing a single leader per round. We believe that a technique similar to the one shown in this work could be applied to bring our fast path to Simplex too. In (Abraham et al., 2021a) Abraham et al. prove that $2\Delta$ is the upper and lower bound for the good-case latency of rotating leader BFT protocols, in the synchronous setting. In this work, we provide a matching upper bound in the partially synchronous setting.
DAG-based BFT protocols. Another line of work proposes to improve the throughput of leader-based protocols by allowing all parties to broadcast transactions simultaneously. Many so-called DAG-based protocols have emerged recently (Keidar et al., 2021; Spiegelman et al., 2022; Babel et al., 2023; Malkhi et al., 2023).
One main idea is to disconnect transactions broadcasting from finalization (Danezis et al., 2022), and another is to allow the finalization of blocks outside the main blockchain by causally referencing them (Lewenberg et al., 2015; Keidar et al., 2021). In Bullshark (Spiegelman et al., 2022), all replicas are parallel leaders and broadcast one batch of transactions in each round using Byzantine Consistent Broadcast. While doing so, they include references to at least $n-f$ blocks from the previous round. A more recent protocol called BBCA-Chain (Malkhi et al., 2023) reduces Bullshark’s latency by removing the need for specialized block layers and instead broadcasts blocks using a primitive called Byzantine Broadcast with Complete-Adopt (BBCA). Mysticeti (Babel et al., 2023) forgoes the confirmation of blocks and instead relies solely on the DAG edges as input to the protocol, leading to better optimistic latency. ICC does not allow causally referencing non-leader blocks to finalize them and thus does not benefit from the same throughput advantages. However, we believe ICC is a good fit to describe our fast path mechanism, and hypothesize that it applies to DAG protocols too.
In Table 1 an overview of state-of-the-art protocols is shown. For both the block finalization latency and the block creation latency, we have included the required number of replicas that have to respond in order to proceed in the optimistic case (synchronous round with a block proposed by an honest leader). While irrelevant in the calculation of the theoretical latency bounds, we found that small changes in these requirements could have large effects in practice, especially on global-scale deployments as they are seen today. (Intuitively, it is not uncommon for a few outlier replicas to have a higher latency. If progress can be made safely without them, a large performance increase can be experienced.)
@@
Other aspects. Many works such as (Malkhi and Nayak, 2023; Camenisch et al., 2022; Gueta et al., 2019) also focus on the message complexity of the fast path and the view change. We do describe how BLS signature aggregation (Boneh et al., 2018) can be used for Banyan, but leave further advancements to future work. Message complexity and performance do not always go hand in hand in practice (Danezis et al., 2022; Camenisch et al., 2022) and oftentimes probabilistic broadcast through gossiping is preferred (Protocol Labs, 2024).